OpenStudy (anonymous):

It costs Rs. 5,000 to paint the total surface of a cube. But in summer, each side expands by 20%. What will be the cost of painting the cube in the summer?

OpenStudy (anonymous):

surface area of the cube=6 a^2 so a is side 6*120/100*120/100a^2=216/25*a^2

OpenStudy (anonymous):

150-216/25a^2

OpenStudy (dumbcow):

looks good so far now lets find the cost we know that 6a^2 = 5000 --> (1.2)^2 (6a^2) = new cost --> (1.2)^2 *5000 = new cost

OpenStudy (anonymous):

7200

OpenStudy (dumbcow):

yep